Let ( , ) ρ = w1 w2
be a portfolio of two securities. Find the value of w1
and w2
in the following
situations:
i) 1 ρ12 = − and ρ is risk-free.
ii) σ1 = σ2
and variance P is minimum.
iii) Variance on P is minimum and 2 ,5.0 ρ12 = − σ1 = and 3 σ2 = .
1
Expert's answer
2020-06-11T17:35:31-0400

i. ρ12 = − and ρ is risk-free.

Rp = w1r1 + w2r2

Variance of the portfolio return

σ2p = w1σ21 + w2σ22 + 2w1w2σ12

i) ρ12 = − and ρ is risk-free.

Value of w1, w2

w1r1 = Rp - w2r2

w1 = Rp /r1 - w2r2/r1

w2 = Rp /r2 – w1r1/r2

ii) σ1 = σ2 and variance P is minimum.

σ = √(w1σ21 + w2σ22 + 2w1w2σ12)

σ2 = √[(w1σ21 + w2σ22 + 2w1w2σ12)]2

σ2 = w1σ21 + w2σ22 + 2w1w2σ12

w1σ21 + 2w1w2σ12 = σ2 - w2σ22  

w121 + 2w2σ12) = σ2 - w2σ22  

w1= (σ21- w2σ22)/ (σ21 + 2w2σ12)

w2= (σ22- w1σ22)/ (σ22 + 2w1σ12)

The answer is the same for different three situations because the portfolio with the two securities hence the minimum variance frontier (MVF) for the three different scenarios portrays a perfect negative correlation (ρ12 = −).